OpenStudy (anonymous):

Use the elimination method to solve the system of equations. Choose the correct ordered pair. 6x + 5y = 61 -5x + 5y = -5 A. (6, 3) B. (6, 5) C. (8, 3) D. (8, 5)

OpenStudy (anonymous):

@ShadowLegendX could you help me?

OpenStudy (shadowlegendx):

Hey @Tallan

OpenStudy (anonymous):

Could you help me out with this problem??

OpenStudy (shadowlegendx):

Ok, do you remember how I taught you substitution a couple days ago? This is similar

OpenStudy (anonymous):

Yes, I remember it a little bit.

OpenStudy (shadowlegendx):

6x + 5y = 61 -5x + 5y = -5 Multiply 6x + 5y = 61 by 5 Multiply -5x + 5y = -5 by 6 30x + 25y = 305 -30x + 30y = -30 Eliminate 25y = 305 30y = -30 Add 55y = 275 Divide both sides by 5 y = 5

OpenStudy (shadowlegendx):

Now this next step is similar to what we did in that other systems of equations that we did. What do you think the next step is @Tallan ?

OpenStudy (anonymous):

Divide?

OpenStudy (shadowlegendx):

Umm no. We already solved for y y = 5

OpenStudy (anonymous):

ohh right sorry then we divide for 6x and -5x?

OpenStudy (shadowlegendx):

Divide? Why would we divide for 6x and -5x o-o

OpenStudy (anonymous):

idk..

OpenStudy (shadowlegendx):

@Tallan I really want you to get this concept so you will be able to do it on your own. 6x + 5y = 61 -5x + 5y = -5 When solving for a system of equations, we are looking for for two points, for the two variables x and y. We just solved for that variable y, but what next after that?

OpenStudy (anonymous):

would we add?

OpenStudy (shadowlegendx):

Why would we add?

OpenStudy (shadowlegendx):

@Tallan I'm trying to understand your thought process here so I can teach you better

OpenStudy (anonymous):

cause you want to get rid of the negative?

OpenStudy (shadowlegendx):

I am not asking for the next step as in, we're going to divide or add. But what is our next purpose in this system of equations. We already solved for y, but what else is there to be solved for?

OpenStudy (anonymous):

solve for x

OpenStudy (shadowlegendx):

YES!!!!!!!!!!!!!!!!!!

OpenStudy (shadowlegendx):

Once you solve for one variable in a system of equations, you look to the other variable, which in this case is x

OpenStudy (anonymous):

Would we have to multiply to find that answer to?

OpenStudy (shadowlegendx):

Now, we are going to try solve for x. Since you are new and just learning the concept, I'll choose the easier one of the two equations for this. -5x + 5y = -5 What do we do with y?

OpenStudy (shadowlegendx):

Remember, y = 5

OpenStudy (anonymous):

5*5

OpenStudy (anonymous):

-5x+5*5=-5

OpenStudy (shadowlegendx):

You're correct, but I want you to think of it this way. You're inputting in 5 for y. "Substituting" That's why I find elimination funny as half of it is just using substitution xD

OpenStudy (anonymous):

haha

OpenStudy (shadowlegendx):

But yes, you're correct. -5x +5(5) = -5

OpenStudy (shadowlegendx):

I want you to try solve for that

OpenStudy (anonymous):

Am I starting it right? -5x +5(5) = -5 -5x +20=-5

OpenStudy (anonymous):

Would it be B? @ShadowLegendX

OpenStudy (shadowlegendx):

Incorrect, 5 x 5 does not equal 25...

OpenStudy (anonymous):

darn it i knew that i just pressed wrong key... I knew that . 5x5 = 25

OpenStudy (shadowlegendx):

@Tallan Are you able to multiply?

OpenStudy (shadowlegendx):

5 x 5 = 5 + 5 + 5 + 5 + 5 Five 5's

OpenStudy (anonymous):

yes/..

OpenStudy (anonymous):

So B is incorrect?

OpenStudy (shadowlegendx):

Well, how did you get B in the first place?

OpenStudy (anonymous):

25/4..... to get 6.25 so i rounded down to 6.. I'm just trying to find the answers quickly.

OpenStudy (shadowlegendx):

I have no idea why you divided 25 by 4. Care to explain that part?

OpenStudy (anonymous):

i am not quite sure. i just divided. xD

OpenStudy (anonymous):

i honestly can't think straight.

OpenStudy (shadowlegendx):

@Tallan When solving for a system of equations, we want to find the x and y points. We want to end up with something like y = 8 and x = 2 To do that, we have to solve/isolate for the variable. Let me do a practice problem 3x = 15 Solve for x

OpenStudy (shadowlegendx):

@Tallan Solve 3x = 15

OpenStudy (anonymous):

it would be 5

OpenStudy (anonymous):

3x5=15

OpenStudy (shadowlegendx):

But what step would you first do to get 5?

OpenStudy (anonymous):

divide 3 to get x alone and divide 3 by 15

OpenStudy (shadowlegendx):

EXACTLY!!! We want to get the variable alone

OpenStudy (shadowlegendx):

But to do that, we have to know when to divide, and divide by the correct numbers. Where you grabbed 4 earlier, I don't know, but lets continue

OpenStudy (anonymous):

would x = 1?

OpenStudy (shadowlegendx):

Nope

OpenStudy (anonymous):

darn...

OpenStudy (shadowlegendx):

Lets start off where we left -5x +5(5) = -5 I believe you corrected your mistake of 5(5) not being 20, but being 25 So that means... -5x + 25 = -5 What would you do now?

OpenStudy (anonymous):

subtract 25 from both sides?

OpenStudy (shadowlegendx):

Yes!!!

OpenStudy (shadowlegendx):

Show me what you get

OpenStudy (anonymous):

-5x=-30

OpenStudy (shadowlegendx):

Correct, so what would you do now?

OpenStudy (anonymous):

-5/-5 -30/-5?

OpenStudy (shadowlegendx):

Yes, divide -5x by 5, and divide -30 by 5 What do you get now?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

so it would be B

OpenStudy (shadowlegendx):

Btw, remember to keep x in your equation, next to -5

OpenStudy (shadowlegendx):

Tallan, that's correct :)

OpenStudy (shadowlegendx):

Now, remember when I taught you substitution, what did I tell you to do after you solved for both points?

OpenStudy (anonymous):

check to make sure its right.

OpenStudy (shadowlegendx):

Yep I want you to input both points into both equations to see if they come out correct :)

OpenStudy (anonymous):

if you take 6 * 6 + 5y = 61 it comes out y=5

OpenStudy (anonymous):

so it is right.

OpenStudy (shadowlegendx):

6(6) + 5(5) = 61 36 + 25 = 61 61 = 61 Correct

OpenStudy (shadowlegendx):

Check the other one Points: y = 5 x = 6

OpenStudy (anonymous):

it comes out 4 -5*5+5y=-5 it came out 4

OpenStudy (anonymous):

y=4?

OpenStudy (shadowlegendx):

-5x + 5y = -5

OpenStudy (shadowlegendx):

How did you get 4? Show me your work ^-^

OpenStudy (anonymous):

nvm.. -5*5+5y=-5 it comes out to -20

OpenStudy (anonymous):

i think i typed it wrong on the calcultor.

OpenStudy (shadowlegendx):

This is the same mistake as last time.

OpenStudy (shadowlegendx):

You think that 5 x 5 is 20

OpenStudy (shadowlegendx):

But it's not...It's 25

OpenStudy (anonymous):

I know, but when i put the equation in my calcultor it came out -20.

OpenStudy (shadowlegendx):

Your calculator is incorrect then -5(6) + 5(5) = -5 -30 + 25 = -5 -5 = -5

OpenStudy (shadowlegendx):

Btw I just noticed that you inputted in the wrong points for the wrong variables

OpenStudy (anonymous):

oh i put that in wrong... thats why i got the wrong answer/. but (6,5) Works..

OpenStudy (shadowlegendx):

Yes @Tallan B is correct. But I want you to get this concept down, as it is just basic algebra and there are much harder concepts

OpenStudy (anonymous):

Believe i am sure there are.... I have 130 to do by the 23. i have to do 17 a day to get everything done on time.

OpenStudy (shadowlegendx):

Alright @ Tallan I'll be on here pretty much all day as today is a free day for me. Tag me when you need help but I may be afk sometimes, so if I can't make it, tag some of my friends like: @camerondoherty @sammixboo @iambatman @pooja195

OpenStudy (anonymous):

Ok Just before i close this question do i have this in the right order? Place the four steps of the elimination method in the proper order. Find the variable with equal or opposite coefficients. If equal, subtract the equations. If opposite, add the equations. Solve the resulting one-variable equation. Substitute the known value into the other equation and solve for the remaining variable.

OpenStudy (shadowlegendx):

Looks good

OpenStudy (anonymous):

Sweet thanks!

OpenStudy (anonymous):

3x - 4y = -30 3x + 2y = 6 would you still multiply for this equation?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

nvm i figured it out.

OpenStudy (shadowlegendx):

You would multiply with the y section